Find the number that is 3/4 of the way from 3/5 to 2 2/3.
Answer by stanbon(75887)

You need to go to the starting point (3/5)
the add (3/4)(difference between 2 2/3 and 3/5)
Find 3/5 + (3/4)(2 2/3 - 3/5)
= 3/5 + (3/4)(8/3 - 3/5)
=3/5 + (3/4)(31/15)
=3/5 +31/20
=12/20 + 31/20
=43/20
Cheers,
Stan H.